# Scrutinize What We Ignore: Reining Task Representation Shift In Context-Based Offline Meta Reinforcement Learning

### v1

@@ 4.2 Monotonic Performance Improvement Concerning Task Representation Shift
Denote $J^{1}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the task encoder and the policy. Similarly, denote $J^{2}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the task encoder and the policy. The performance difference bound in the COMRL setting can be defined as
@@
when $C$ is non-negative, the algorithm allows a monotonic performance improvement in the real environment.
According to Definition 4.4, we need to find the positive $C$ to improve the performance monotonically. To achieve this, we transform Eq. (7) to the following formulation, where $J^{mutual}(\theta_{i}),i\in[1,2]$ denote the expected return of the policy $\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{i})$ among the expectation of tasks.
$J^{2}(\theta_{2})-J^{1}(\theta_{1})$ $=(J^{2}(\theta_{2})-J^{mutual}(\theta_{2}))+(J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1}))$
$+(J^{mutual}(\theta_{1})-J^{2}(\theta_{1}))+(J^{2}(\theta_{1})-J^{1}(\theta_{1}))$ (8)
Then, we introduce a mild assumption for the task representation distribution.
@@
By leveraging Eq. (4.2) and the above assumptions, we derive the lower bound of the performance difference.
@@
For the need of monotonic performance improvement guarantee, we further let the right-hand side of Eq. (10) be positive. Then, we would have the following requirement to achieve monotonicity.
@@
The above equation unveils the key influence of the task representation shift $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ for two aspects. If the task representation shift is too small, $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ may greater than the right-hand side of Eq. (11). On the contrary, if the task representation shift is too large, the right-hand side of Eq.(11) may be negative. Both of these cases would lead to performance improvement collapse. Thus, we argue that the impacts of task representation shift cannot be ignored.
To give an example for this problem, we introduce the following assumptions.
@@
The i.i.d samples from $Z_{mutual}(\cdot|x)$ can be obtained.
Assumption 4.10.
The task encoder is trained by fitting an empirical distribution on i.i.d samples from $Z_{mutual}(\cdot|x)$.
Though sampling from $Z_{mutual}(\cdot|x)$ seems impractical, we can approximate it as sampling from a task encoder pre-trained by various bounds of $I(Z;M)$.
Corollary 4.11 (Monotonic performance improvement guarantee on empirical distribution).
Denote $\kappa$ as $\frac{2R_{max}}{(1-\gamma)^{2}}$. Given that the task encoder has already been trained by $N$ i.i.d samples from $Z_{mutual}(\cdot|x)$. When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8, Assumption 4.9 and Assumption 4.10, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the task encoder from extra $k$ i.i.d samples, where:
$k=\frac{8\kappa^{2}L_{z}^{2}}{(\epsilon^{mutual}_{12}-\kappa L_{z}\beta)^{2}}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}-N$ (12)
@@ 4.3 Practical Implementation
To simplest make it practical, we first transform Eq. (11) to an optimization problem.
$\min|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|,\,\,s.t.\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|>0$ (13)
As alluded in Section 4.1, many contrastive learning-based algorithms involved in prior works [li2020focal, yuan2022robust, gao2024context, li2024towards] can be adopted to minimize $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$. Unfortunately, no one among them is free from parameter-tuning. We propose to use a classifier-based manner to address this concern. Specifically, after extracting from the task encoder, the task representation would be passed through a fully connected layer to get the probabilities for each task. Then, the task encoder is trained by the supervision of the task label and back-propagation of the cross-entropy loss. Note that this classifier-based manner enjoys tuning-free property and also can be seen as an approximation of $I(Z;M)$. To the best of our knowledge, we are the first to propose adopting a classifier-based manner to train the task encoder in the COMRL setting.
For reining task representation shift $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$, since the values lying in the constraint of Eq. (13) are intractable, approximately estimating them may introduce a series of hyper-parameters thus significantly increasing the complexity of the algorithm. Instead, one insight from Corollary 4.11 is that tuning the update of the task encoder may bring better performance improvements. Through the experiment, we find that adjusting the update frequency of the task encoder to half of the origin is sufficient to beat the existing SOTA algorithms. Empirical results demonstrate that the previously ignored task representation shift does make a remarkable difference.
Our main algorithm is shown in Algorithm 1. Compared to the backbone, RETRO only adds one line of code, which is colored by red. To maintain the consistency of prior COMRL works, we also adopt BRAC [wu2019behavior] to train the policy. Moreover, combining prior works with our approach to reining task representation shift is also ablated in the experiments.
Algorithm 1 RETRO Framework
0:  Offline Datasets $\bm{X}=\{x^{i}\}_{i=1}^{N_{train}}$, training tasks $\bm{M}=\{m^{i}\}_{i=1}^{N_{train}}$, initialized learned policy $\pi_{\theta}$, Q function $Q_{\psi}$, context encoder $q_{\phi}$, task batch size $N_{tbatch}$
1:  while not done do
2:     Sample task $\{m^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{M}$ and the corresponding replay buffer =$\mathcal{R}=\{x^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{X}$
3:     for step in each iteration do
4:        Sample context $\{\bm{c}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
5:        Obtain task representations $\{\bm{z}^{j}\sim q_{\phi}(\bm{z}|\bm{c}^{j})\}_{j=1}^{N_{tbatch}}$
6:        // Rein the task representation shift
7:        if step % 2 == 0 then
8:           Compute $\mathcal{L}_{\text{RETRO}}=\textit{CrossEntropyLoss}(z^{j},m^{j})$
9:           Update $\phi$ to minimize $\mathcal{L}_{\text{RETRO}}$
10:        end if
11:        Detach the task embeddings $\{\bm{z}^{j}\}_{j=1}^{N_{tbatch}}$
12:        Sample training data $\{\bm{d}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
13:        Compute $\mathcal{L}_{\text{BRAC}}$ to update $\theta,\psi$
14:     end for
15:  end while
@@ 5 Experiment
Our experiments are conducted to answer 4 questions: 1) Compared to existing SOTA algorithms, can RETRO achieve more stable performance improvements and higher asymptotic performance? 2) Can RETRO be integrated into prior works like UNICORN and FOCAL? 3) How does reining the task representation shift with different frequencies affect RETRO? 4) Compared to existing SOTA algorithms, can RETRO achieve lower training-time consumption?
@@ 5.1 Environments Settings
We adopt MuJoCo [todorov2012mujoco] and MetaWorld [yu2020meta] benchmarks to evaluate the algorithms. We choose 4 competitive baselines to validate our superiority. The baselines cover UNICORN [li2024towards], which adopts reconstruction loss to balance the impacts of spurious correlation, CSRO [gao2024context], which adopts CLUB loss to upper-bound the impacts of the behavior policy, FOCAL [li2020focal], which directly applies distance metric learning to train the task encoder. and Classifier, which is our backbone without reining the task representation shift. The implementation details of all algorithms are shown in Appendix Section 7.3. Note that all these baselines can be categorized into the return discrepancy scheme stated in Section 4.1. Following the protocol of UNICORN, we randomly sample 20 training tasks and 20 testing tasks from the task distribution. We train the SAC [haarnoja2018soft] agent from scratch for each task and use the collected replay buffer as the offline dataset. During the meta-testing phase, we adopt a fully offline setting in a few-shot manner like [li2020focal] that randomly samples a trajectory from the dataset as the context to obtain the task representation and then passes it to the meta-policy to complete the testing in the true environments.
5.2 Baseline Comparisons
@@
Figure 1: Testing returns of RETRO against baselines on six benchmarks. Solid curves refer to the mean performance of trials over 4 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 1 illustrates the learning curves of RETRO against other baselines on 20 testing tasks for each benchmark. The results showcase that RETRO can unanimously achieve SOTA asymptotic performance across all benchmarks. Compared to the backbone Classifier, with adding only one line of code to rein the task representation shift, RETRO achieves remarkable performance improvements (e.g. approximately 20% relative performance improvements on Ant-Dir and 25% relative performance improvements on Dial-Turn). It is worth noting that RETRO shows comparably stable performance improvements in all the benchmarks especially on Point-Robot and Dial-Turn. Recall that RETRO only needs to adjust the update frequency to half of its origin, thus leading to a tuning-free algorithm. On the contrary, both UNICORN and CSRO are especially sensitive to the hyper-parameter settings. This phenomenon is particularly evident in the extremely challenging benchmark Humanoid-Dir, which holds the dimension of state space 376. We argue that without carefully parameter-tuning trials, UNICORN and CSRO cannot exhibit the desired performance. Admittedly, though more advanced algorithms may exist to better control the task representation shift, adopting the simplest manner to achieve SOTA performance makes RETRO more applicable to other fields.
5.3 Can RETRO Be Integrated Into Other Methods?
According to our analysis in Section 4.1, prior COMRL works ignore the impacts of task representation shift, which may lead to performance collapse. As our main focus is to rein the task representation shift, which is orthogonal to the prior works, RETRO should bring better performance improvements when being integrated into prior works. We denote RETRO-UNICORN and RETRO-FOCAL as combining the reining method into the original UNICORN and FOCAL respectively. To ensure fairness, we also adjust the update frequency of the task encoder in RETRO-UNICORN and RETRO-FOCAL to half of the origin.
@@ 5.3 Can RETRO Be Integrated Into Other Methods?
Figure 2: Testing returns of RETRO-UNICORN and RETRO-FOCAL against the origin on three benchmarks. Solid curves refer to the mean performance of trials over 4 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 2 demonstrates the learning curves of RETRO-UNICORN and RETRO-FOCAL against their backbones on 20 testing tasks for each benchmark. The results illustrate that adopting RETRO enables better performance improvements across all the benchmarks, thereby validating our core claim on reining the impacts of task representation shift. Note that RETRO(-Classifier) has comparable or even higher performance than RETRO-UNICORN, we believe that RETRO(-Classifier) should be a preferred option due to the tuning-free property.
Though our goal is to design the simplest algorithm to prove the effectiveness of reining the task representation shift, we can also notice that the convergence speed is relatively lower than the backbones at the beginning of the training process. Hence, in the future, we envision designing a strong algorithm to control the task representation shift delicately, where the springboard is RETRO.
5.4 The Influence Of Different Update Frequencies
@@
Figure 3: Testing returns of difference update frequency on two benchmarks. Solid curves refer to the mean performance of trials over 4 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 4: Average training time consumption of RETRO against other baselines on six benchmarks.
Figure 4 illustrates the influence of different update frequencies on the performance of RETRO. The results show that setting the update frequency to half of the origin can generally strike a balance among the convergence speed, training stability and asymptotic performance across all the benchmarks. Hence, we argue that adopting RETRO as a tuning-free algorithm is acceptable.
5.5 Can RETRO Reduce Training-Time Consumption?
We report the training-time consumption of RETRO against other baselines across all benchmarks in Figure 4. All algorithms are trained with the same architecture on two RTX3090 graphic cards. With the exception of Humanoid-Dir, which can only run one seed per graphic card due to the limitation, the reported times for the remaining benchmarks pertain to running two seeds per card and a total of four seeds concurrently during training. Figure 4 shows that RETRO has a significant time-consumption reduction against UNICORN [li2024towards], CSRO [gao2024context] and FOCAL [li2020focal]. Though this fact is based on our choice of backbone to be a classifier, it is the one proposed for the first time in our paper aiming at achieving a tuning-free algorithm. Compared to the backbone classifier, RETRO also achieves lower training-time consumption, as RETRO reduces the update frequency of the task encoder.

### v2

@@ 4.2 Monotonic Performance Improvement Concerning Task Representation Shift
Denote $J^{1}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the context encoder and the policy. Similarly, denote $J^{2}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the context encoder and the policy. The performance difference bound in the COMRL setting can be defined as
@@
when $C$ is non-negative, the algorithm allows a monotonic performance improvement.
According to Definition 4.4, we need to find the positive $C$ to improve the performance monotonically. We denote $J^{mutual}(\theta_{i}),i\in[1,2]$ as the expected return of the policy $\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{i})$ among the expectation of tasks and then introduce a mild assumption for the task representation distribution.
@@
By leveraging Definition 4.4 and the above assumptions, we derive the lower bound of the performance difference.
@@
Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of the performance difference in COMRL can be formalized as:
@@
For the need of the monotonic performance improvement guarantee, we further let the right-hand side of Eq. (9) be positive. Then, we would have the following condition to achieve monotonicity.
@@
The above equation unveils the key influence of the variation of task representation $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ for two aspects. If $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ is too small, $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ may greater than the right-hand side of Eq. (10). On the contrary, if $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ is too large, the right-hand side of Eq.(10) may be negative. Both of these cases would lead to the performance improvement collapse. Thus, we argue that the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ cannot be ignored.
To highlight the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$, we name this issue task representation shift. Under some mild assumptions, we can conclude that it is possible to achieve monotonic performance improvement on the condition of proper context encoder updates, as shown in Theorem 4.10.
@@
There exists an $\alpha$ for any given $Z_{1}(\cdot|x)$, $|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|^{2}\leq\frac{\alpha}{b}$, where $Z_{2}(\cdot|x)$ denotes the context encoder updated by maximizing $I(Z;M)$ with the data size $b$ from the training dataset based on $Z_{1}(\cdot|x)$ and $Z^{empirical}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $b$ i.i.d samples from $Z_{mutual}(\cdot|x)$ based on $Z_{1}(\cdot|x)$  [^2] We set discrepancy to be subject to an inverse proportional function is because as the training samples increase, $Z_{2}$ and $Z^{empirical}_{2}$ should become similar..
Theorem 4.10 (Monotonic performance improvement guarantee on training process).
Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex. Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent. When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8 and Assumption 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder with maximizing $I(Z;M)$ from at least extra $k$ samples, where:
$k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (11)
@@
Theorem 4.10 shows the connection between the needed update data size $k$ and the performance improvement brought only by the policy update $\epsilon^{mutual}_{12}$. If the calculated $k$ is larger than the given batch size, the context encoder should avoid this update and wait for the accumulation of $\epsilon^{mutual}_{12}$. As $\epsilon^{mutual}_{12}$ increases, $k$ decreases. When $k$ is smaller than the given batch size, the monotonic performance improvement can be achieved by updating the context encoder. Notice that this updating process can be performed multiple times, as long as satisfying Assumption 4.7. The updating process is visualized in Figure 1 right. Compared to the original alternating framework, Theorem 4.10 unveils that the core part for better performance is to adjust the update of the context encoder to rein in the task representation shift, thereby showcasing the advantage over the previous works. The general algorithmic framework is shown in Algorithm 1, where the way to adjust the update of the context encoder is colored by red.
Algorithm 1 General Algorithmic Framework Towards Reining In The Task Representation Shift
0:  Offline training datasets $\bm{X}$, initialized policy $\pi_{\theta}$, context encoder $q_{\phi}$ and given task representation shift threshold $\beta$
1:  for iter in alternating iterations do
2:     // Update the context encoder
3:     if Conditions for updating the context encoder are met then
4:        while Accumulated task representation shift is less than $\beta$ do
5:           Sample context from $\bm{X}$
6:           Obtain task representations from $q_{\phi}$ with inputting the sampled context
7:           Compute $\mathcal{L}_{\text{encoder}}$ concerning maximizing $I(Z;M)$
8:           Update $\phi$ to minimize $\mathcal{L}_{\text{encoder}}$
9:        end while
10:     end if
11:     // Update the policy
12:     Detach the task representations
13:     Sample training data from $\bm{X}$
14:     Compute $\mathcal{L}_{\text{policy}}$ via standard offline RL algorithms
15:     Update $\theta$ to minimize $\mathcal{L}_{\text{policy}}$
16:  end for
@@ 4.3 Practical Implementation
To show the impacts of the task representation shift issue, we set different manners to adjust the context encoder. We denote the given batch size as $bs$ and the accumulation steps of task representation shift as $acc$. Our manners cover 1) $k=2\times bs,acc=1$, 2) $k=3\times bs,acc=1$, 3) $k=1\times bs,acc=2$, and 4) $k=1\times bs,acc=3$. All manners as well as the original manner, namely $k=1\times bs,acc=1$ are performed for 8 different random seeds.
With respect to the specific context encoder learning algorithms concerning maximizing $I(Z;M)$, we choose 3 representative algorithms that are also widely adopted in previous COMRL endeavors. The reconstruction-based algorithm, passing the state, action, and task representation through the decoder to reconstruct the next state and the reward signal, is applied in [zintgraf2019varibad, dorfman2021offline, li2024towards] and is proven to be the lower bound of $I(Z;M)$. The contrastive-based algorithm, using the task representation to compute distance metric learning loss, is applied in [li2020focal, gao2024context, li2024towards] and is proven to be the upper bound of $I(Z;M)$ as well as contain the lower bound of $I(Z;M)$. The cross-entropy-based algorithm is a straightforward discrete approximation to $I(Z;M)$ and is proposed by our work. Specifically, after extracting from the context encoder, the task representation would be passed through a fully connected layer to get the probabilities for each task. Then, the context encoder is trained by the supervision of the task label and back-propagation of the cross-entropy loss. The implementation details of these algorithms are shown in Appendix Section 8.3.
With respect to the policy learning algorithm, to maintain the consistency of previous COMRL works, we directly adopt BRAC [wu2019behavior] to train the policy.
@@ 5 Experiment
Our experiments are conducted to show that the previous optimization framework that ignores the impacts of task representation shift is not sufficient. We hope to illustrate the necessity of reining in the task representation shift, laying the foundation for further research towards better asymptotic performance of COMRL.
@@ 5.1 Environments Settings
We adopt MuJoCo [todorov2012mujoco] and MetaWorld [yu2020meta] benchmarks to evaluate the algorithms. Following the protocol of UNICORN  [li2024towards], we randomly sample 20 training tasks and 20 testing tasks from the task distribution. We train the SAC [haarnoja2018soft] agent from scratch for each task and use the collected replay buffer as the offline dataset. During the meta-testing phase, we adopt a fully offline setting in a few-shot manner like [li2020focal] that randomly samples a trajectory from the dataset as the context to obtain the task representation and then passes it to the meta-policy to complete the testing in the true environments.
5.2 MAIN RESULTS
@@
Figure 2: Testing returns of different manners to rein in the task representation shift against the original manner concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 2 illustrates the learning curves of our different manners to rein in the task representation shift against the original manner on 20 testing tasks for each benchmark concerning 3 widely adopted objectives for maximizing $I(Z;M)$. We find that these specific objectives hold unanimous performance demonstration on all benchmarks that the previous training framework, namely $k=1\times bs,acc=1$, cannot provide the best asymptotic performance. These objectives focus solely on the optimization target of the task representation, namely the optimal solution of $I(Z;M)$. While they are guaranteed to produce reasonable results, as shown in Section 4.1, ignoring the impacts of task representation shift still hinders the capability of these algorithms to reach their full potential. As shown by Figure 2, even with minor changes to the algorithms, the performance improvements are substantial.
@@
Figure 3: Testing returns of different manners to rein in the task representation shift against the original manner concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$ on different data qualities of Ant-Dir. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
5.3 Can The Results Show Consistency Across Different Data Qualities?
To make our claim more universal, we conduct the experiment on different data qualities. We collect three types of datasets with size being equal to the dataset used in Section 5.2. Depending on the quality of the behavior policies, we denote these datasets as random, medium, and expert respectively.
Figure 3 demonstrates the learning curves of our different manners to rein in the task representation shift against the original manner concerning 3 widely adopted objectives for maximizing $I(Z;M)$ on Ant-Dir. We find that the contrastive-based algorithm fails on the random dataset. Except for this case, the others show consistent results that reining in the task representation shift may lead to better performance improvements.
Based on these impressive results, we believe that developing a smarter algorithm to control the task representation shift automatically is an appealing direction and we leave this to future work.
6 Discussion
This section seeks to raise some interesting issues derived from our work. We hope these issues can further enable pondering on the significance of the task representation shift.
@@ 6 Discussion
Figure 4: Testing returns of the pre-training scheme against training from scratch of cross-entropy, contrastive, and reconstruction based objectives on Ant-Dir. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
6.1 Can The Pretraining Scheme Be Adopted To Achieve Better Performance Improvement?
According to our analysis, the task representation shift issue only happens in the case that the context encoder needs to be trained from scratch. Hence, a question is raised naturally. If we train the context encoder in advance and use this pre-trained context encoder to train the policy directly, can we achieve better performance improvement?
To answer this question, we conduct the experiment on Ant-Dir and still use the cross-entropy, contrastive, and reconstruction based objectives to pre-train the context encoder. As shown in Figure 4, there is a significant performance gap between pre-training and training from scratch. To better theoretically interpret this, we introduce the following Corollary.
Corollary 6.1 (Monotonic performance improvement condition for pre-training scheme).
Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, Assumption 4.5, the monotonic performance improvement condition for pre-training scheme is:
$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (12)
Proof: See Section 8.2 in the Appendix. ∎
According to Corollary 6.1, pre-training cannot actually achieve monotonic performance improvements as $|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|$ is not equal to 0. As the task representation cannot be changed under the pre-training scheme, compared to training from scratch, this scheme loses some degrees of freedom to improve performance.
Additionally, this scheme can be intuitively regarded as degrading to the problem that how to design a pre-training algorithm to better approximate $I(Z;M)$. Since the cross-entropy based objective is the direct approximation of $I(Z;M)$, holding better performance under the pre-training scheme is also in line with expectation.
This analysis further enhances the importance of considering the impacts stemming from the variation of the task representation in the alternating process. Nevertheless, how to achieve better performance under the pre-training scheme is also an interesting problem.
Figure 5: The 2D projection of the learned task representation space in Ant-Dir. Points are uniformly sampled from the evaluation datasets. Tasks of given goals from 0 to 6 are mapped to rainbow colors, ranging from purple to red.
6.2 Can The Visualization Of Task Representation Be Strongly Relied Upon to Imply The Asymptotic Performance?
Previous COMRL endeavors [li2020focal, yuan2022robust, gao2024context] mostly apply t-SNE on the learned task representation to show the differentiation of the tasks. They hold the belief that better performance corresponds to better differentiation of the tasks. However, based on our analysis, the visualization result only depends on the task representation at convergence and ignores the task representation variation during the whole optimization process. According to the illustration shown in Figure 5, less desirable differentiation results can also lead to better performance. Hence, it is insufficient to imply performance based on such evaluation principles. Nevertheless, we recognize that the visualization result can be seen as an auxiliary metric to assist in determining the task representation.